OpenStudy (anonymous):

GIVING MEDALS & FANNING . The equation of line EF is y = x + 6. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0, −2) y = −2x − 2 y = x + 2 y = x − 2 y = − 2x + 2

OpenStudy (mathstudent55):

The equation of line EF is in slope-intercept form. Can you tell what the slope of line EF is?

OpenStudy (anonymous):

See if this helps any. y = mx + b (Slope-Intercept Form) m = slope b = y-intercept y = x + 6 (Your equation) m = ? b = ?